In a certain Algebra 2 class of 28 students, 14 of them play basketball and 7 of them play baseball. There are 2 students who play both sports. What is the probability that a student chosen randomly from the class plays basketball or baseball?




1
Expert's answer
2022-05-06T11:36:01-0400

"2" students in the class play both sports. It means that "14-2=12" students play only basketball and "7-2=5" students play only baseball. Thus, there are "2+12+5=19" students that play basketball or baseball. The probability that a randomly chosen student plays basketball or baseball is "\frac{19}{28}\approx0.6786". (it is rounded to "4" decimal places)

Answer: "0.6786". (it is rounded to "4" decimal places)